# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v5

@@ 1 Introduction
Preference optimization has become the dominant paradigm for aligning large language models with human values, with Direct Preference Optimization (DPO) [1] emerging as a simple and effective alternative to traditional reinforcement learning from human feedback (RLHF) [2, 3]. However, DPO’s reliance on hard, pairwise preference labels makes it brittle. In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy-tailed label noise, DPO’s performance degrades by up to 71% relative to our method (0.472 vs. 0.809 WinMass). This sensitivity arises because the DPO objective only regularizes differences of log-probabilities, providing a weak inductive bias that is susceptible to overfitting corrupted labels.
We introduce Anchored Direct Preference Optimization (ADPO), a unified framework that connects seemingly disparate learning paradigms. Through an anchored projection—minimizing $\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}\big)$ where $\tilde{p}_{\theta}$ is defined by anchored logits $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all arise from different choices of the target distribution $q$, the anchor policy $\pi_{\mathrm{ref}}$, and the temperature $\tau_{\mathrm{anc}}$ (Section 3, Proposition 3.2). By centering the coordinate system at a reference policy before applying softmax, ADPO provides groupwise shift invariance and Fisher-metric regularization—properties absent in standard methods.
@@
Collectively, our work makes four key contributions: (1) we propose ADPO as a unified projection framework that subsumes major learning paradigms (SFT, KD, max-ent RL, DPO) and robustly handles soft and listwise data; (2) we establish the theoretical connection between anchoring and implicit, Fisher-metric trust regions; (3) we uncover the task-dependent trade-off between anchor strategies, a key practical insight; and (4) we provide extensive empirical evidence and clear guidance for selecting anchor strategies in different learning contexts, informing both online RLHF and offline distillation pipelines.
@@ 2 Related Work
Preference optimization. More generally, the Preference Optimization (PO) framework [18] formulates an entropy-regularized objective aligning policy probabilities with qualitative preference signals, which inspires our entropy-regularized interpretation of anchoring (Section 3), though our work focuses on preference optimization for RLHF-style alignment rather than combinatorial optimization. DPO [1] reparameterizes the reward model through policy-reference log-ratios, enabling direct optimization. Extensions include identity-free preferences (IPO [5]), contrastive objectives (CPO [6]), and iterative refinement [7]. Our work provides a unified framework encompassing these variants.
@@
Conventional knowledge distillation (KD) [22] aligns a student with a teacher by minimizing $\mathrm{KL}(q\|p_{\theta})$, which directly matches absolute probabilities in the (unanchored) probability simplex. While effective, this objective lacks a fixed geometric center and thus provides no notion of a trust region, making it sensitive to teacher noise and distributional shifts. In contrast, trust-region policy optimization (TRPO) [23] explicitly constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, defining a moving local quadratic region that stabilizes updates. Our ADPO formulation can be viewed as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, the optimization is performed in coordinates centered at the reference, while the local quadratic well (Fisher metric) is centered at the target distribution $q$, yielding an implicit trust region without an explicit KL constraint. We are not aware of prior work that formulates an anchored cross-entropy maintaining a fixed geometric center in probability space.
@@
the optimal policy admits the Boltzmann form [19, 21]
@@
so any relative quantity (pairwise/listwise) cancels $Z(s)$. Eqs. (1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ to these soft targets. A full derivation of (1) as the minimizer of $\mathbb{E}_{s}\,\mathrm{KL}\big(\pi(\cdot|s)\,\|\,\pi^{*}(\cdot|s)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [18] (see their Eqs. (3)–(6), (20)–(27)).
@@
For any context-independent shift $h(s)$, if $r^{\prime}(s,a)=r(s,a)-h(s)$ then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance. See [18], Prop. 3.1 and App. D.2.
3 Anchored Direct Preference Optimization
@@
$x$: context (prompt/state); $S_{x}=\{y_{1},\ldots,y_{|S_{x}|}\}$: candidate set (actions/outputs). $\pi_{\theta}(y\mid x)$: student policy; $\pi_{\mathrm{ref}}(y\mid x)$: anchor (fixed or dynamic). $s_{y}=\log\pi_{\theta}(y\mid x)$, $s^{\mathrm{ref}}_{y}=\log\pi_{\mathrm{ref}}(y\mid x)$: log-probabilities. $q(y\mid S_{x})$: target preference distribution (from teacher, rewards, or labels). $u_{y}=(s_{y}-s^{\mathrm{ref}}_{y})/\tau_{\mathrm{anc}}$: anchored logits. $\tilde{p}_{\theta}(\cdot\mid S_{x})=\mathrm{softmax}(u)$: anchored model distribution.
Table 1: Temperature parameters and their roles across ADPO formulations.
Symbol Role Where used
$\beta_{r}$ Teacher/target inverse temperature Forms $q$ from rewards: $q\propto\exp(R/\beta_{r})$
$\beta$ Student inverse temperature (pairwise) Pairwise logistic: $\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))$
$\tau_{\mathrm{anc}}$ Anchoring temperature Defines anchored logits: $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ (listwise)
Remark (scale non-identifiability): In pairwise settings, temperatures $\beta$ and $\beta_{r}$ only appear through their ratio $\beta_{r}/\beta$ in the Bayes-optimal condition, so the statistical optimum is invariant to jointly rescaling them. We set $\beta=\beta_{r}=1.0$ by default. However, the optimization dynamics are affected: pairwise gradients scale as $\beta$ (since $\partial\ell/\partial\theta=\beta(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij})(\nabla s_{i}-\nabla s_{j})$), while listwise gradients scale as $1/\tau_{\mathrm{anc}}$ (because $\partial\mathcal{L}/\partial s_{i}=\tfrac{1}{\tau_{\mathrm{anc}}}(\tilde{p}-q)$). Thus smaller $\beta$ slows pairwise updates; smaller $\tau_{\mathrm{anc}}$ tightens the trust region via curvature $\propto 1/\tau_{\mathrm{anc}}^{2}$ but increases gradient magnitudes.
Remark (temperature $\tau_{\mathrm{anc}}$ and regularization strength): In listwise settings with anchoring, $\tau_{\mathrm{anc}}$ controls the balance between the anchor and target via the closed-form solution $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ (Proposition 3.4). Intuitively:
  • As $\tau_{\mathrm{anc}}\to 0$: $q^{\tau_{\mathrm{anc}}}\to\mathbf{1}$ (uniform), so $\pi_{\theta}\to\pi_{\mathrm{ref}}$ (stronger regularization toward the anchor).
  • As $\tau_{\mathrm{anc}}\to\infty$: $q^{\tau_{\mathrm{anc}}}$ becomes increasingly peaked at $\operatorname*{arg\,max}q$, so $\pi_{\theta}$ tracks the target $q$ more aggressively.
Note that, when the optimum is reachable, $\tilde{p}_{\theta}$ matches $q$ at optimality for any $\tau_{\mathrm{anc}}$; the temperature only controls how the underlying policy relates to the anchor via $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$. Moreover, locally around the optimum, a smaller $\tau_{\mathrm{anc}}$ induces a tighter implicit trust region: deviations in $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ are penalized more strongly since $\|\nabla_{s}\mathcal{L}\|\propto 1/\tau_{\mathrm{anc}}$ and the local curvature scales as $1/\tau_{\mathrm{anc}}^{2}$.
3.1 Core Formulation: Anchored Distribution as the Protagonist
ADPO performs preference learning via anchored projection: we define a distribution $\tilde{p}_{\theta}$ in coordinates centered at a reference policy $\pi_{\mathrm{ref}}$, then minimize its KL divergence to a target $q$. All theory revolves around this anchored distribution $\tilde{p}_{\theta}$.
Definition of anchored distribution.
Given student policy $\pi_{\theta}$, reference anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}>0$, we define:
  • Anchored logits:
    $u_{i}\;=\;\frac{\log\pi_{\theta}(i\mid x)-\log\pi_{\mathrm{ref}}(i\mid x)}{\tau_{\mathrm{anc}}}.$ (3)
  • Anchored distribution:
    $\tilde{p}_{\theta}(i\mid S_{x})\;=\;\mathrm{softmax}(u)_{i}\;=\;\frac{\exp(u_{i})}{\sum_{j\in S_{x}}\exp(u_{j})}.$ (4)
  • Objective: Given target distribution $q(\cdot\mid S_{x})$,
    $\min_{\theta}\ \mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\Big[\mathrm{KL}\!\big(q(\cdot\mid S_{x})\,\|\,\tilde{p}_{\theta}(\cdot\mid S_{x})\big)\Big].$ (5)
Remark 3.1 (Sequence-level scoring).
For autoregressive models, we use length-normalized (per-token average) log-probabilities for $s$ and $s^{\mathrm{ref}}$ to avoid length bias. This preserves softmax affine-invariance, so all theoretical results (shift invariance, Fisher geometry, closed-form solutions) apply to the sequence-level distribution over candidates.
@@ 3.1 Core Formulation: Anchored Distribution as the Protagonist
Since $\tilde{p}_{\theta}$ depends only on log-odds differences $u_{i}-u_{j}=(s_{i}-s_{i}^{\mathrm{ref}})-(s_{j}-s_{j}^{\mathrm{ref}})$ (via softmax), adding any constant $c$ to all $s_{i}$ and $s_{i}^{\mathrm{ref}}$ within a group leaves $\tilde{p}_{\theta}$ unchanged (Lemma 2.1). This cancels group-level biases—e.g., annotator-specific baselines—automatically.
Closed-form optimal solution.
When the anchored distribution $\tilde{p}_{\theta}$ can be matched exactly to the target $q$ (e.g., sufficient expressivity, full action coverage), the optimal policy takes the multiplicative fusion form (Proposition 3.4):
$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau_{\mathrm{anc}}}.$ (6)
Proof sketch: $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=q$ iff $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}=\log q+c\mathbf{1}$, i.e., $\pi_{\theta}\propto\pi_{\mathrm{ref}}q^{\tau_{\mathrm{anc}}}$. ∎
Fisher geometry = implicit trust region.
The KL objective $\mathrm{KL}(q\|\tilde{p}_{\theta})$ admits a second-order expansion around the optimum $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$):
$\mathrm{KL}(q\|\tilde{p}_{\theta})\;\approx\;\tfrac{1}{2}\,\delta^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\delta\;=\;\tfrac{1}{2\tau_{\mathrm{anc}}^{2}}\,\operatorname{Var}_{q}\!\big[s-s^{\mathrm{ref}}-\tau_{\mathrm{anc}}\log q\big],$ (7)
where $\delta=u-u^{\star}$ is $q$-centered (Lemma 3.5). The matrix $\operatorname{Diag}(q)-qq^{\top}$ is the softmax Fisher metric, inducing a local quadratic trust region whose curvature depends only on $q$ (not on the anchor $\pi_{\mathrm{ref}}$). This provides a distribution-space analogue of TRPO’s trust region, but implicit—no explicit KL constraint needed.
DPO as a binary special case.
When $|S_{x}|=2$, the anchored distribution reduces to $\tilde{p}_{\theta}(i)=\sigma((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}})$ where $\Delta_{\theta}=s_{i}-s_{j}$ and $\Delta_{\mathrm{ref}}=s_{i}^{\mathrm{ref}}-s_{j}^{\mathrm{ref}}$. With Bradley–Terry targets $q(i)=\sigma(\beta_{r}\Delta R)$ and uniform $\pi_{\mathrm{ref}}$ (so $\Delta_{\mathrm{ref}}=0$), Eq. (5) recovers the standard DPO loss [1]. For listwise $|S_{x}|>2$, ADPO generalizes to Plackett–Luce targets while retaining the Fisher-metric trust region.
Unification of learning paradigms.
Different choices of $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$ recover major learning objectives as special cases:
Proposition 3.2 (ADPO unifies major learning paradigms).
Let $\mathcal{L}_{\mathrm{ADPO}}$ be defined in Eq. (5). Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}$ recover major learning objectives as special cases:
  1. 1.
    Supervised fine-tuning (SFT). When $q(\cdot\mid S_{x})$ is the data label distribution, $\pi_{\mathrm{ref}}$ is uniform, and $\tau_{\mathrm{anc}}=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, recovering standard cross-entropy loss.
  2. 2.
[…]
  3. 3.
    Maximum-entropy RL. When $q(a\mid s)\propto\exp(Q(s,a)/\alpha)$ and $\pi_{\mathrm{ref}}$ is uniform, setting $\tau_{\mathrm{anc}}=1$ recovers the soft-optimal Boltzmann policy $\pi_{\theta}\propto\exp(Q/\alpha)$ [19, 21].
  4. 4.

### v6

@@ 1 Introduction
Preference optimization has become the dominant paradigm for aligning large language models with human values, with Direct Preference Optimization (DPO) [1] emerging as a simple and effective alternative to traditional reinforcement learning from human feedback (RLHF) [2, 3]. However, DPO’s reliance on hard, pairwise preference labels makes it brittle. In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy noise, DPO’s performance degrades by up to 93% relative to our method (0.375 vs. 0.725 WinMass). This sensitivity arises because the DPO objective only regularizes differences of log-probabilities, providing a weak inductive bias that is susceptible to overfitting corrupted labels.
We introduce Anchored Direct Preference Optimization (ADPO), a unified framework that connects seemingly disparate learning paradigms. Through an anchored projection—minimizing $\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}\big)$ where $\tilde{p}_{\theta}$ is defined by anchored logits $(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all arise from different choices of the target distribution $q$, the anchor policy $\pi_{\mathrm{ref}}$, and the temperature $\tau_{\mathrm{anc}}$ (Section 3.2, Proposition 3.4). By centering the coordinate system at a reference policy before applying softmax, ADPO provides groupwise shift invariance and Fisher-metric regularization—properties absent in standard methods.
@@
[…]
@@ 2 Related Work
Preference optimization. More generally, the Preference Optimization (PO) framework [18] formulates an entropy-regularized objective aligning policy probabilities with qualitative preference signals, which inspires our entropy-regularized interpretation of anchoring (Section 3.2), though our work focuses on preference optimization for RLHF-style alignment rather than combinatorial optimization. DPO [1] reparameterizes the reward model through policy-reference log-ratios, enabling direct optimization. Extensions include identity-free preferences (IPO [5]), contrastive objectives (CPO [6]), and iterative refinement [7]. Our work provides a unified framework encompassing these variants.
@@
Conventional knowledge distillation (KD) [22] aligns a student with a teacher by minimizing $\mathrm{KL}(q\|p_{\theta})$, which directly matches absolute probabilities in the (unanchored) probability simplex. While effective, this objective lacks a fixed geometric center and thus provides no notion of a trust region, making it sensitive to teacher noise and distributional shifts. In contrast, trust-region policy optimization (TRPO) [23] explicitly constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, defining a moving local quadratic region that stabilizes updates.
Our ADPO formulation can be viewed as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, the optimization is performed in coordinates centered at the reference, while the local quadratic well (Fisher metric) is centered at the target distribution $q$, yielding an implicit trust region without an explicit KL constraint. To our knowledge, a direct anchored cross-entropy formulation emphasizing a fixed geometric center in probability space has not been formalized in prior work, though related ideas appear in advantage-weighted methods [24, 25] and maximum a posteriori policy optimization [26].
@@
the optimal policy admits the Boltzmann form [19] (also known as energy-based policies [21])
@@
so any relative quantity (pairwise/listwise) cancels $Z(s)$. Eqs. (1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ to these soft targets. A full derivation of (1) as the minimizer of $\mathbb{E}_{s}\,\mathrm{KL}\big(\pi(\cdot|s)\,\|\,\pi^{*}(\cdot|s)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [18] (see their Eqs. (3)–(6), (20)–(27)).
@@
For any context-independent shift $h(s)$, if $r^{\prime}(s,a)=r(s,a)-h(s)$ then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(\ell-\ell^{\mathrm{ref}})$—inherits groupwise shift invariance.
Proof sketch.
For any group-constant $c$, $\mathrm{softmax}(u)=\mathrm{softmax}(u+c\mathbf{1})$ by shift-invariance. Since $r^{\prime}(s,a)=r(s,a)-h(s)$ only changes the group constant (log-partition $\log Z(s)$ in Eq. (2)), the Bradley–Terry/Plackett–Luce probabilities and anchored distribution $\tilde{p}$ remain unchanged. See [18], Prop. 3.1 and App. D.2 for full derivation. ∎
3 Anchored Direct Preference Optimization (ADPO)
3.1 Problem Setting and Anchored Softmax
Motivation: Why anchoring?
Traditional preference optimization methods (e.g., DPO with uniform reference) operate on absolute log-probabilities $\ell_{i}=\log\pi_{\theta}(y_{i}|x)$, making them sensitive to:
  • Annotator baseline biases: Different annotators may have systematically different baseline preferences
  • Sequence-level biases: Length bias in language models, state-value offsets in RL
  • Distribution shift: Changes in the reference distribution between training and deployment
ADPO addresses these issues by introducing a reference anchor $\pi_{\mathrm{ref}}$ and working with relative log-probabilities $(\ell-\ell^{\mathrm{ref}})$, which automatically cancels group-level baseline shifts through the softmax’s shift-invariance property (detailed below).
@@ 3.1 Problem Setting and Anchored Softmax
Let $x\!\sim\!\mathcal{D}$ be a context (e.g., prompt or state). For a candidate set $S_{x}=\{y_{1},\dots,y_{P}\}$ (e.g., responses or actions), we define:
  • Student policy $\pi_{\theta}(y|x)$ with log-probabilities $\ell_{i}=\log\pi_{\theta}(y_{i}|x)$
  • Anchor policy $\pi_{\mathrm{ref}}(y|x)$ with log-probabilities $\ell_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(y_{i}|x)$
  • Target preference $q(\cdot|S_{x})$ with $q\succ 0$ (strictly positive distribution)
We work with anchored logits:
$u_{i}\;=\;\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\,,\qquad\bar{u}_{i}\;=\;u_{i}-\sum_{j\in S_{x}}q(j|S_{x})\,u_{j},$ (3)
where $\tau_{\mathrm{anc}}>0$ is an anchoring temperature controlling regularization strength, and $\bar{u}$ denotes the $q$-centered anchored logits (always applied for numerical stability and geometric consistency).
@@
Because $\tilde{p}_{\theta}$ depends only on differences of anchored logits $u_{i}-u_{j}$, adding a constant to all $(\ell_{i}-\ell_{i}^{\mathrm{ref}})$ within a group leaves $\tilde{p}_{\theta}$ unchanged. This property automatically cancels group-level baselines such as:
  • Annotator-specific biases (different annotators may have different baseline preferences)
  • Sequence-level biases (e.g., length bias in language models)
  • State-dependent baselines in RL (e.g., value function offsets)
Remark 3.1 (Sequence tasks and length normalization).
For sequence generation tasks, one typically constructs $\ell$ and $\ell^{\mathrm{ref}}$ using length-normalized log-probabilities to avoid length bias. The affine invariance of the groupwise softmax ensures all results remain valid.
Target distribution $q$.
The target $q(\cdot|S_{x})$ encodes the desired preference over candidates within group $S_{x}$. When the full action/candidate space is considered, we write $q(\cdot|x)$; the in-group distribution can be viewed as conditioning on $S_{x}$. Common choices include:
  • Plackett–Luce (PL) soft labels from rankings
  • Bradley–Terry (BT) soft labels from pairwise comparisons: $q(i\succ j|S_{x})=\sigma(\beta_{r}\Delta R_{ij})$
  • Q-function based: $q(a|s)\propto\exp(Q(s,a)/\beta_{r})$ in RL settings
  • Teacher distribution: $q=\pi_{\text{teacher}}$ for knowledge distillation
3.2 Unified ADPO Objective and Its Special Cases
Our master objective is the groupwise forward KL divergence:
$\boxed{\min_{\theta}\ \mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\Big[\mathrm{KL}\!\big(q(\cdot|S_{x})\,\big\|\,\tilde{p}_{\theta}(\cdot|S_{x})\big)\Big],\quad\tilde{p}_{\theta}=\mathrm{softmax}(\bar{u}),\ \bar{u}=\tfrac{(\ell-\ell^{\mathrm{ref}})}{\tau_{\mathrm{anc}}}-\langle q,\tfrac{(\ell-\ell^{\mathrm{ref}})}{\tau_{\mathrm{anc}}}\rangle\mathbf{1}.}$ (4)
(Equivalently, using $u$ without centering yields the same $\tilde{p}$ up to a constant shift, by softmax shift-invariance.)
Why forward KL?
Forward KL $\mathrm{KL}(q\|\tilde{p}_{\theta})$ is mode-covering (encourages $\tilde{p}_{\theta}$ to place mass wherever $q$ does), making it robust to soft/noisy targets; reverse KL is mode-seeking (concentrates on a single mode), which is brittle under label noise. For listwise supervision with uncertain preferences, mode-covering is more stable.
Within a group, this equals a cross-entropy up to a constant. Writing $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ where $A(u)=\log\sum_{j}e^{u_{j}}$ is the log-partition function, the loss is strictly convex in $u$ (on $\mathbf{1}^{\perp}$):
$\sum_{i}q_{i}\log\tilde{p}_{\theta,i}\;=\;-\frac{1}{\tau_{\mathrm{anc}}}\sum_{i}q_{i}\big(\ell_{i}-\ell_{i}^{\mathrm{ref}}\big)\;+\;\log\!\sum_{j}\exp\!\Big(\frac{\ell_{j}-\ell_{j}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\Big)\;=\;\mathrm{KL}(q\|\tilde{p}_{\theta})+H(q).$ (5)
Closed-form optimum in distribution space.
If the model class is expressive enough, the optimum of (4) satisfies:
Proposition 3.2 (Closed-form optimum).
When $\tilde{p}_{\theta}(\cdot|S_{x})=q(\cdot|S_{x})$ is attainable, the optimal policy satisfies
$\pi_{\theta}(y|x)\;\propto\;\pi_{\mathrm{ref}}(y|x)\,q(y|x)^{\tau_{\mathrm{anc}}}.$ (6)
This requires mild support conditions: in the discrete case, $q(y|x)>0\Rightarrow\pi_{\mathrm{ref}}(y|x)>0$; in the continuous case, absolute continuity and integrability ensure $\tilde{p}=q$ is attainable and the proportionality is normalizable (well-defined). When training on sampled groups $S_{x}$, this gives the conditional optimum within each group; if groups are uniformly sampled from the full action space and the model is sufficiently expressive, the empirical optimum converges to the population optimum.
Proof sketch.
At optimum, $\tilde{p}_{\theta}(i|S_{x})=q(i|S_{x})$. By definition,
$\tilde{p}_{\theta}(i|S_{x})=\mathrm{softmax}\!\Big(\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\Big)\propto\exp\!\Big(\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\Big)=\Big(\frac{\pi_{\theta}(i|x)}{\pi_{\mathrm{ref}}(i|x)}\Big)^{1/\tau_{\mathrm{anc}}}.$
Setting this equal to $q(i|S_{x})$ and solving yields (6). ∎
Temperature interpretation.
  • Small $\tau_{\mathrm{anc}}$: Strong regularization toward anchor $\pi_{\mathrm{ref}}$ (conservative updates). The local curvature scales as $1/\tau_{\mathrm{anc}}^{2}$ (Section 3.3), yielding a tighter implicit trust region.
  • Large $\tau_{\mathrm{anc}}$: Weak regularization, $\pi_{\theta}$ tracks $q$ more closely (aggressive updates)
  • $\tau_{\mathrm{anc}}=1$: Balanced regime (commonly used in practice)
Pairwise and DPO as corollaries.